Find angle between the tangents at t = 1 and t = 2 to the curve ๐‘Ÿฬ… = ๐‘ก2 ๐‘–ฬ…+ (๐‘ก3 โˆ’ 2๐‘ก) ๐‘—ฬ… + (3๐‘ก โˆ’ 4) ๐‘˜ฬ….


1
Expert's answer
2021-05-21T07:26:49-0400

"r'(t)=2ti+(3t^2-2)j+3k"

r'(1)=2i+ 1j+3k

r'(2)=4i+10j+3k

cos(r'(1),r'(2))=(2*4+1*10+3*3)/"(\sqrt{r'(1)}*\sqrt{r'(2)})" =27/"(\sqrt{14}*\sqrt{125})" =27/"(5\sqrt{70})" angle=arcos(27/"(5\sqrt{70}))" )